OpenStudy (anonymous):

Which is the equation of a line that passes through the point (3, -4) and has a slope of 2.

OpenStudy (anonymous):

y = mx + b y = 2x + b -4 = 2(3) + b b = -10 y = 2x - 10

OpenStudy (anonymous):

thanks :D